# MARS: A Multi-Agent Framework Incorporating Socratic Guidance for Automated Prompt Optimization

### v1

@@ 3.2 Main Results
Figure 4: The Efficiency Analysis chart presents a comparison of the PE metric between MARS and other baseline methods across different tasks. For each task, different models are grouped together in a single bar chart cluster, allowing for a clear visual comparison of performance efficiency differences.
Table 2 presents the experimental results of MARS on domain-specific tasks, covering areas like Chinese, law, and mathematics, all of which require specialized knowledge and reasoning. In these tasks, MARS outperforms the previous SOTA methods to 6.42%, further demonstrating its ability to better guide LLMs in domain-specific knowledge discovery and application. This not only lowers the barrier to utilizing LLMs but also enhances their generalization capability. Moreover, compared to the original prompt and CoT(ZS), MARS achieves improvements of 25.31% and 20.72%, respectively, underscoring its effectiveness and practicality in these specialized domains.
3.3 Efficiency Analysis
The balance between resource consumption and performance improvement is a crucial analysis metric (Yang et al., 2024b). In the process of MARS during APO, the phased optimization of prompts gradually offers a great advantage in balancing this trade-off, allowing for performance improvement while maintaining efficient resource consumption. To quantify this balance, we propose a new metric–PE (Prompt Efficiency). The formula is as follows:
$\text{PE}=\frac{Accuracy}{Consumption},$ (6)
where accuracy refers to the performance of the APO tasks, while the consumption refers to the numbers of calling the LLMs API to optimize prompt. Clearly, the higher the PE metric, the better the model performs with less resource consumption, reflecting the efficiency of resource utilization.
As shown in Figure 6, MARS is compared with strong baselines across multiple APO tasks, specifically evaluating its PE metric on four different tasks. The results indicate that in the tasks of Boolean Expression and Formal Fallacies, the PE value of MARS is more than twice that of the other baselines. Meanwhile, in tasks Ruin Name and Sports Understanding, MARS also outperforms other baselines by more than 20%, further demonstrating its superior capability in balancing resource consumption and performance. This further highlights the guiding capability and superiority of the MARS model’s approach, which first conducts task planning and then progressively decomposes and optimizes. By adopting this strategy, MARS can allocate resources more effectively, minimize unnecessary consumption, and ensure both efficiency and stability in the optimization process.
4 Supplementary Analysis
To conduct more comprehensive experiments, this section will cover the following topics: ablation experiments in Section 4.1, convergence analysis in Section 4.2, and efficency analysis in Section 4.3.
Meanwhile, to comprehensively demonstrate the generalization capability of the MARS framework, Appendix C presents the experimental performance of MARS when using GPT-4o as the base model, as well as the applicability and optimization results of MARS’s main experiments across various other LLMs. To explore the impact of sample size on the experimental results, the sample size analysis is shown in Section D. Additionally, Appendix E provides a detailed list of the specific prompts used by each intelligent agent, offering a clearer understanding of their roles and operational logic. To illustrate the collaborative mechanism of the multi-agent system of MARS, Appendix F documents the complete interaction and optimization process for a single APO task. Finally, Appendix G contains the optimized prompts for the 17 APO tasks conducted in the main experiment of this study, supporting further replication and analysis.
4.1 Ablation Study
Variation B.E. D.QA F.F. G.S. R.N. S.U. Avg.
@@ 4.1 Ablation Study
Table 3: Performance under different ablation settings are analyzed. We performed ablation experiments on the planner module ${w/o}_{\text{Plan}}$, the Teacher-Critic-Student module ${w/o}_{\text{Soc}}$, and the Critic Agent ${w/o}_{\text{Cri}}$ to evaluate the impact of removing these components.
Table 3 presents the impact of removing the Planner agent, the Teacher-Critic-Student Socratic guidance dialogue module, and the Critic agent on the overall performance. It can be observed that removing the Teacher-Critic-Student Socratic guidance dialogue module causes the greatest loss in performance, followed by the Planner module, while the removal of the Critic agent has the least impact on the results.
Removing the Teacher-Critic-Student Socratic guidance dialogue module has the greatest impact on performance. Without this module, after the Planner formulates optimization steps, the iterative process to refine the prompt is absent. The system simply sends raw steps to the target agent, reducing optimization effectiveness. Additionally, removing the Planner agent causes a noticeable drop in performance. The Planner structures the optimization steps in the APO task, and without it, the Socratic guidance dialogue module lacks optimization guidance, negatively affecting overall performance. Lastly, removing the Critic agent affects the Socratic guidance Dialogue module, particularly in providing feedback to the Teacher. As shown in Table 3, this results in a 3.55% performance loss, highlighting its smaller but still meaningful role in the optimization process.
4.2 Converagence Analysis
@@
Figure 5 presents the convergence analysis across four BBH tasks. To better monitor the APO process, we visualized its iterative optimization trajectory, selecting 10 iterations as the observation window.
@@
The results show that MARS exhibits an upward training trajectory in the early iterations. For instance, in Task A, MARS converges to the optimal solution by iteration 5. In contrast, in the OPRO task, even after 10 iterations, convergence is not achieved, leading to higher resource consumption. This highlights MARS’s efficiency in reaching the optimal solution within fewer iterations, thereby reducing resource overhead and improving optimization efficiency.
@@
4.3 Case Study and Interpretability
This section focuses on the interpretability analysis of Figure 3 and Figure 6. As mentioned earlier, the MARS model provides both process interpretability and result interpretability.
Process interpretability is mainly reflected in two aspects: First, the planning steps are fully text-based, explicitly outlining the optimization path guidance, such as the planned steps in Figure 6. Second, the interpretability of the Socratic iterative process. Taking the Teacher-Critic-Student iterative process in Figure 3 as an example, the Teacher asks Socratic-style questions, the Critic follows the “assess-adjust” process, and finally, the Student iterates step by step. The entire process is completely transparent and has practical significance, reflecting the iterative process of each substep.
Result interpretability is evident in the results shown in Figure 6, such as the requirements for fields like dynamic tolerance threshold, validation and refinement, etc. These are specific requirements for the task, allowing for clear understanding and guiding the agents to achieve optimal performance.
Figure 6: Case study of a geometry shape task, showing the planning step and new prompt after optimized iterations.
5 Related Works
Prompt Optimization. Early work primarily focused on two aspects: discrete optimization of hard prompts (Shin et al., 2020; Wen et al., 2024; Chen et al., 2023; Zhang et al., 2022) and continuous vector optimization of soft prompts (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2024b). However, these methods are highly task-dependent and exhibit locality. With the advent of LLMs, traditional methods have become outdated. APE (Zhou et al., 2022) pioneered the use of generative methods to optimize instructions. Since APE, there have been two major approaches. The first approach (Zhou et al., 2022; Xu et al., 2023; Pryzant et al., 2023; Wang et al., 2023) is the generate-search model, where multiple candidate sequences are generated, and methods like Monte Carlo search are used to optimize the prompt. The second approach (Yang et al., 2024a; Ye et al., 2023) is the meta prompts method, where sophisticated meta prompts are designed to optimize the prompt. In contrast to these two approaches, MARS employs a planned optimization path, iteratively generating high-quality prompts. This approach alleviates the inefficient search in prompt spaces issues in the first approach and addresses the challenges of limited flexibility of fixed templates in the second approach.
@@ 5 Related Works
Multi-Agent. Based on LLMs, a combination of AI agents capable of performing specific functions forms a multi-agent system  (Richards, 2023; Yang et al., 2023; Wu et al., 2023). Given a statement of a specific task, AI agents can attempt to break complex problem statements into subtasks and use tools, including data retrieval from the internet, to solve them step-by-step through automatic iterations. Some studies (Poldrack et al., 2023; Wang et al., 2024a; Xi et al., 2025; Ni and Gao, 2021) use multi-agent systems to address issues such as problem identification, code development and debugging, plotting results and analysis, and providing interactive feedback with the human user. Ni and Buehler (2024) demonstrates the potential of organizing an AI multi-agent collaborative team to automatically solve mechanical problems, showcasing an enhanced ability to understand, formulate, and validate engineering problem solutions through self-correction and mutual correction. Inspired by their work, we leverage multi-agent technology to autonomously plan the APO optimization path and design a Teacher-Critic-Student collaborative approach for iterative optimization.
6 Conclusion
This study introduces a MARS method. Specifically, we develop a multi-agent framework incorporating Socratic guidance that includes a Planner agent and six task-specific agents for APO. First, to tackle the issue of limited flexibility of fixed templates, we utilize the Planner agent to autonomously design optimization paths for various tasks, ensuring that each task adheres to its own specific optimization trajectory. Next, we employ a Teacher-Critic-Student Socratic guidance dialogue pattern to iteratively refine the prompts while address the issue of inefficient search in prompt spaces. Finally, we validate the optimized prompts within the Target agent and iterate until the optimal prompt is identified. We conduct extensive experiments across a range of general tasks and domain-specific datasets to assess the effectiveness of MARS and explore the interpretability of the optimization process and results.
Limitations
MARS represents the first attempt to apply a multi-agent architecture to APO tasks, achieving improvements. Notably, the planner LLM’s autonomous optimization path planning and the Teacher-Critic-Student Socratic-style iterative prompt optimization have shown good results. However, we acknowledge two limitations: (1) the question of whether there exists a more universal representation of prompts across different task types, highlighting the need for broader prompt design patterns that can accommodate a variety of tasks; and (2) the consideration of incorporating environmental feedback into the APO process to enhance the system’s interactivity and error correction capabilities. We aim to address these issues in the future.
Appendix A Tasks and Datasets
@@
General Task Evaluation
We select six tasks from the BBH (Suzgun et al., 2022) and MMLU (Wang et al., 2024b) datasets, respectively. BBH tasks consist of six challenging reasoning tasks that assess logical inference and problem-solving skills, including boolean expressions, disambiguation QA, formal fallacies, geometric shapes, ruin names, and sports understanding. MMLU tasks cover six subject-specific tasks designed to evaluate general knowledge across diverse fields, including college biology, college medicine, electrical engineering, high school world history, human aging, and marketing.
Domain-Specific Task Evaluation
We include three benchmarks: C-Eval (Huang et al., 2024), GSM8K (Zhang et al., 2024a), and LSAT-AR (Zhong et al., 2023). C-Eval is a Chinese evaluation benchmark that covers domain-specific topics such as art studies, clinical medicine, and Urban and Rural Planner. GSM8K is a widely used mathematical reasoning dataset. LSAT-AR focuses on legal reasoning, evaluating AI performance in law-related tasks.
Dataset Split
@@
Appendix B Experiment Settings and Baselines
We compare MARS with three categories of baselines: original prompts, CoT prompts, and some of the latest APO methods. Specifically, (1) original prompts refer to the prompts used in the datasets, where each dataset often provides some initial guidance for the tasks. (2) To build the CoT (Zero-Shot) baseline, we add the prompt Let’s think step by step at the beginning of each task; based on this, we further include a specific example to create the CoT (Few-Shot) baseline. (3) Finally, we compare MARS with some strong baseline methods from recent years, including Automatic Prompt Engineer (APE) (Zhou et al., 2022), Prompt Optimization with Textual Gradients (ProTeGi) (Pryzant et al., 2023), Optimization by PROmpting (OPRO) (Yang et al., 2024a), and Prompt Engineer 2 (PE2) (Ye et al., 2023). APE and ProTeGi generate multiple prompts and perform search optimization to find the optimal prompt, while OPRO and PE2 optimize prompts by designing a sophisticated meta prompts.
We select a powerful LLM, deepseek-V2.5-1210 (Guo et al., 2025), as our primary agent for the APO tasks. Not only does deepseek-V2.5-1210 exhibit strong reasoning and generation capabilities in a variety of natural language processing tasks, but it also efficiently explores multiple angles when facing complex prompt optimization requirements, making it well-suited for adapting to different tasks and datasets in the APO process. We adopt accuracy as our primary evaluation metric to comprehensively assess the performance of different methods across various task scenarios. Table 1 presents our experimental results on 12 general task datasets, illustrating the performance of APO in diverse scenarios, while Table 2 summarizes its performance on five domain-specific datasets, underscoring the model’s versatility and stability across different fields. To further validate the generality and robustness of our method, we additionally employed another high-performance LLM, GPT-4o (Achiam et al., 2023), for extended comparative experiments, with the corresponding findings reported in Appendix C. Furthermore, we provide a detailed description of our experimental setup and procedures in the appendix to facilitate future research and practical applications.
Appendix C Generalization Across Different Base and Target Models
@@
C.1 Base Model of GPT-4o
To verify the generality and effectiveness of the proposed method in this study, we conduct further experiments by replacing the base model with GPT-4o (Achiam et al., 2023).
@@ C.1 Base Model of GPT-4o
C.2 Other Target LLMs
We also conduct testing of the main experimental results of MARS on other Target LLMs, using the most optimized prompts from this study to validate the method across multiple different LLMs. These tests include performance on Deepseek-R1, GPT-3.5, GPT-4, and GPT-4o. The results in Table 6 from these experiments demonstrate that the optimization achieved on the Deepseek-V2.5 base model possesses strong generalization ability. Even on other powerful LLMs, MARS continues to show excellent performance and optimization effects. This further validates the cross-model applicability of the MARS method, proving that it is not limited to a single base model but can achieve impressive results across various advanced LLM platforms.
@@ C.2 Other Target LLMs
Base Deepseek GPT Avg.
-V2.5 -R1 -3.5 -4 -4o
Origin 56.96 61.48 44.79 49.70 55.84 53.75
CoT(ZS) 62.72 73.82 63.45 66.94 70.38 67.46
MARS 79.59 83.05 69.30 73.21 80.86 77.20
Appendix D Sample Size Analysis
One of the key highlights of this study is that the training data consists of only a single sample—we utilize just one data point from the current task for training. To analyze the rationality of one-shot training, we present a comparison of 0-shot, 1-shot, 3-shot, and baseline methods in Table 7.
Tasks Train B.E. D.QA F.F. G.S. R.N. S.U.
@@ Appendix D Sample Size Analysis
Table 7: Performance comparison of different sampling strategies on the evaluation metric. Train means the trining data.
@@
Appendix E Prompts for Agents
Table 8 summarizes the prompts used for all agents in this paper, with each agent playing a crucial role in the overall process. First, the Manager is responsible for managing and invoking all agents. As the coordinator of the entire process, it ensures that each agent operates in the correct sequence and manner. Next, the UserProxy receives the specific task details and forwards them to the subsequent stages of the process. The role of the UserProxy is to convert external inputs into a format that the system can handle, ensuring that task details are transmitted accurately.
In the following stages, the Planner creates a structured plan based on the task requirements, providing a clear path for execution with defined steps and objectives. The Planner’s role is not just to map out the overall flow of the task but also to guide the Teacher, Student, and Critic through iterative refinements, enhancing the quality of the solution. Throughout this process, the Teacher provides guidance, the Student makes adjustments and learns, and the Critic evaluates and gives feedback. Finally, the Target validates the results to ensure their accuracy and effectiveness, guaranteeing that the output meets the desired objectives. This validation process is a critical step in ensuring the reliability and effectiveness of the entire system.
Appendix F Full-process Prompt Optimization
Figure 7 presents a detailed example of the full-process prompt optimization, demonstrated using the Geometry Shapes task from the BBH dataset. The figure clearly outlines the overall optimization framework and its key stages, effectively illustrating the relationships and operations at each step.
From the Manager, UserProxy, and Planner to the iterative refinement within the Teacher-Critic-Student framework, each phase is visually represented. The figure sequentially presents the functions of each agent and their collaboration, showcasing how optimization and coordination are effectively carried out at every stage, ensuring the accuracy and effectiveness of the final result.
Appendix G Universal Optimum Solution

### v2

@@ Introduction
Figure 5: Inference-time scaling law. The horizontal axis denotes the inference-time computational cost, while the vertical axis represents the average performances on all tasks.
MARS achieves strong and consistent performance gains across domain-specific tasks, highlighting its effectiveness in knowledge-intensive reasoning. Table 2 presents the experimental results of MARS on domain-specific tasks, covering areas like Chinese, law, and mathematics, all of which require specialized knowledge and reasoning. In these tasks, MARS outperforms the previous SOTA methods to 6.42%, further demonstrating its ability to better guide LLMs in domain-specific knowledge discovery and application. This not only lowers the barrier to utilizing LLMs but also enhances their generalization capability. Moreover, compared to the original prompt and CoT(ZS), MARS achieves improvements of 25.31% and 20.72%, respectively, underscoring its effectiveness and practicality in these specialized domains.
Efficiency Analysis
MARS Consistently Achieves the Highest Computational Efficiency.
The balance between resource consumption and performance improvement is a crucial analysis metric (yang2024harnessing). As shown in Figure 5, MARS consistently outperforms all baseline methods in terms of computational efficiency, as demonstrated by its superior inference-time scaling behavior across multiple APO tasks.
Notably, under the same number of output tokens, MARS achieves the highest performance across all evaluated tasks. Conversely, to reach comparable performance levels, baseline methods require more output tokens—indicating higher computational cost.
These results highlight MARS’s strong ability to balance performance and resource usage through its structured optimization strategy. By performing high-level task planning followed by step-wise Socratic refinement, MARS enables more efficient resource allocation, reduces unnecessary computation, and ensures both effectiveness and robustness throughout the APO process.
Variation B.E. D.QA $F.$ G.S. R.N. S.U. Avg.
@@
Table 3: Performance under different ablation settings are analyzed. We performed ablation experiments on the planner module ${w/o}_{\text{Plan}}$, the Teacher-Critic-Student module ${w/o}_{\text{Soc}}$, and the Critic Agent ${w/o}_{\text{Cri}}$ to evaluate the impact of removing these components. w/o indicates the experiment was run without the specified module.
Supplementary Analysis
To further validate the effectiveness of MARS, we conduct three additional analyses in this section: an ablation study to assess the contribution of each component, a convergence analysis to examine the optimization stability over iterations, and an investigation of the sensitivity to Other Target LLMs.
In addition, Appendix D reports MARS’s generalization performance when applied to GPT-4o and different initial prompt $p_{0}$. Appendix E analyzes the effect of sample size on performance. Appendix F provides the internal prompts used by each agent to clarify their roles. Appendix G presents the full multi-agent interaction process on a representative APO task. Appendix H offers the optimized prompts for all 17 tasks to facilitate reproducibility.
Ablation Study
The Socratic dialogue mechanism plays the most critical role in MARS, as shown by the largest performance drop upon its removal. Table 3 presents the impact of removing three key components: the Planner agent, the Teacher-Critic-Student Socratic module, and the Critic agent. Removing the Socratic module leads to the most substantial degradation, as the system loses its iterative refinement capability and sends unprocessed sub-goals directly to the Target agent, resulting in poor optimization quality. Eliminating the Planner also causes a notable drop, since the Socratic dialogue lacks structured guidance without its sub-goal trajectory. Finally, while the Critic contributes less overall, its feedback loop with the Teacher improves prompt quality; removing it leads to a 3.55% performance loss, as shown in Table 3.
Converagence Analysis
@@
MARS achieves faster convergence in most tasks, improving both efficiency and optimization quality. Figure 6 presents the convergence analysis across four BBH tasks. To better monitor the APO process, we visualize the iterative optimization trajectory within a 10-iteration observation window.
@@
The results show that MARS exhibits an upward reward trend in the early stages. For instance, in Task ‘Ruin Names’, it converges to the optimal solution by iteration 5. In contrast, in the OPRO task, convergence is not reached even after 10 iterations, resulting in higher resource consumption. This comparison highlights MARS’s ability to reach optimal prompts in fewer steps, reducing computational cost and enhancing efficiency.
Base Deepseek GPT Avg.
-V2.5 -R1 -3.5 -4 -4o
Origin 56.96 61.48 44.79 49.70 55.84 53.75
CoT(ZS) 62.72 73.82 63.45 66.94 70.38 67.46
MARS 79.59 83.05 69.30 73.21 80.86 77.20
Other Target LLMs
MARS demonstrates strong cross-model generalization, maintaining high performance across diverse LLM backbones. We further evaluate MARS on additional Target LLMs—Deepseek-R1, GPT-3.5, GPT-4, and GPT-4o—using the optimized prompts from previous experiments to assess robustness across model families. As shown in Table 4, prompts optimized on the Deepseek-V2.5 base model generalize well, preserving strong performance even on larger or structurally different LLMs. MARS consistently achieves notable gains across models, validating its model-agnostic design and broad applicability.
Related Works
Prompt Optimization.
Early work primarily focused on two aspects: discrete optimization of hard prompts (shin2020autoprompt; wen2024hard; chen2023instructzero; zhang2022tempera) and continuous vector optimization of soft prompts (lester2021power; li2021prefix; liu2024gpt). However, these methods are highly task-dependent and exhibit locality. With the advent of LLMs, traditional methods have become outdated. APE (zhoularge) pioneered the use of generative methods to optimize instructions. Since APE, there have been two major approaches. The first approach (zhoularge; xureprompting; pryzant2023automatic; wang2023promptagent) is the generate-search model, where multiple candidate sequences are generated, and methods like Monte Carlo search are used to optimize the prompt. The second approach (Yang0LLLZC24; ye2023prompt; zhang2025mars) is the meta prompts method, where sophisticated meta prompts are designed to optimize the prompt. In contrast to these two approaches, MARS employs a planned optimization path, iteratively generating high-quality prompts. This approach alleviates the inefficient search in prompt spaces issues in the first approach and addresses the challenges of limited flexibility of fixed templates in the second approach.
@@
Multi-Agent.
@@
Based on LLMs, a combination of AI agents capable of performing specific functions forms a multi-agent system  (richards2023auto; yang2023auto; wu2023autogen; zhang2025gkg). Given a statement of a specific task, AI agents can attempt to break complex problem statements into subtasks and use tools, including data retrieval from the internet, to solve them step-by-step through automatic iterations. Some studies (poldrack2023ai; wang2024survey; xi2025rise; ni2021deep; lin2025foundation) use multi-agent systems to address issues such as problem identification, code development and debugging, plotting results and analysis, and providing interactive feedback with the human user. ni2024mechagents (zhang2025maps) demonstrates the potential of organizing an AI multi-agent collaborative team to automatically solve mechanical problems, showcasing an enhanced ability to understand, formulate, and validate engineering problem solutions through self-correction and mutual correction. Inspired by their work, we leverage multi-agent technology to autonomously plan the APO optimization path and design a Teacher-Critic-Student collaborative approach for iterative optimization.
Conclusion
We propose MARS, a novel multi-agent framework for adaptive APO that integrates Socratic guidance within a POMDP formulation. It includes: (1) a Planner that generates task-specific optimization trajectories, and (2) a Teacher-Critic-Student dialogue enabling interpretable prompt refinement. This simulates pseudo-gradient paths in discrete prompt space, narrowing the search scope. Modeled as a POMDP: the Student’s latent state is the hidden state, Teacher-Critic interactions define actions, and prompt outputs serve as observations. A Target agent guides iteration via performance rewards. Experiments show MARS consistently outperforms baselines while maintaining transparent optimization trajectories.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62137002, 62277042, 62293553, 62450005, 62437002, 62477036, 62477037, 62176209, 62192781, 62306229), the “LENOVO-XJTU” Intelligent Industry Joint Laboratory Project, the Shaanxi Provincial Social Science Foundation Project (No. 2024P041), the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-593), and the Youth Innovation Team of Shaanxi Universities “Multi-modal Data Mining and Fusion”.
Proof of Proposition
Proof of Proposition 1 (Socratic Policy Improvement Bound)
Let $\Pi=\{\pi_{\text{t}},\pi_{\text{c}},\pi_{\text{s}}\}$ denote the joint policy composed of the Teacher, Critic, and Student agents. Let $p_{i}$ denote the prompt state at refinement step $i$, and define the associated latent state as $s_{i}$. At each step, the action $a_{i}=(q_{i},c_{i})$ induces a transition from $s_{i-1}$ to $s_{i}$, and yields a prompt $p_{i}=\pi_{\text{s}}(a_{i},p_{i-1})$.
We assume the existence of an underlying task reward function:
$\mathcal{R}(p)=\mathbb{E}_{(x,y)\sim D}\left[f\left(\pi_{\text{tar}}(x;p),y\right)\right],$ (11)
which is Lipschitz-continuous with constant $\lambda>0$. Our goal is to lower-bound the cumulative reward improvement from $p_{0}$ to $p_{n}$ based on Socratic signal quality.
Step 1: Define Local Advantage.
Let the local reward gain be defined as:
$A_{i}:=\mathcal{R}(p_{i})-\mathcal{R}(p_{i-1}).$ (12)
We assume the expected advantage at each step satisfies:
$\mathbb{E}[A_{i}]\geq\bar{A}_{i}>0,$ (13)
i.e., the expected contribution of the composite action $a_{i}$ is beneficial.
Step 2: Control for Socratic Variance.
We define the conditional variance of $A_{i}$ given the state $s_{i-1}$ as:
$\text{Var}[A_{i}\mid s_{i-1}]\leq\sigma^{2}.$ (14)
This variance captures uncertainty due to imperfect questions or noisy feedback in the Socratic interaction.
Step 3: Apply Jensen-Bernstein Inequality.
Since $\mathcal{R}(p)$ is Lipschitz, and prompt updates occur in a discrete space, we use a Bernstein-style bound:
$\mathbb{E}[\mathcal{R}(p_{i})]\geq\mathcal{R}(p_{i-1})+\bar{A}_{i}-\frac{\sigma^{2}}{2\lambda}.$ (15)
Step 4: Accumulate Across Steps.
Summing over all $n$ refinement steps, we obtain:
$\mathbb{E}[\mathcal{R}(p_{n})]-\mathcal{R}(p_{0})$ $=\sum_{i=1}^{n}\mathbb{E}[A_{i}]$ (16)
$\geq\sum_{i=1}^{n}\left(\bar{A}_{i}-\frac{\sigma^{2}}{2\lambda}\right).$
This completes the proof of the lower bound. It shows that as long as the Socratic actions are informative ($\bar{A}_{i}>0$) and the variance $\sigma^{2}$ is controlled, the cumulative reward is guaranteed to improve linearly with the number of refinement steps.
Proof of Proposition 2 (Monotonic Reward Stability)
Let $\{p_{i}\}_{i=0}^{n}$ denote the sequence of prompts generated by the MARS refinement process, where each prompt is updated via:
$p_{i}=\pi_{\text{s}}(a_{i},p_{i-1}),\quad\text{with}\quad a_{i}=(q_{i},c_{i}),$ (17)
i.e., a Socratic action $a_{i}$ is used to refine the prompt at step $i$.
Assume the reward function $\mathcal{R}:\mathcal{P}\rightarrow\mathbb{R}$ is $\lambda$-Lipschitz continuous with respect to the prompt representation, i.e., for all $i=1,\dots,n$,
$|\mathcal{R}(p_{i})-\mathcal{R}(p_{i-1})|\leq\lambda\cdot\|p_{i}-p_{i-1}\|,$ (18)
where $\|\cdot\|$ denotes a norm over the prompt space (e.g., token-level edit distance or embedding-based distance).
Suppose further that the update step size is bounded as:
$\|p_{i}-p_{i-1}\|\leq\varepsilon.$ (19)
Then, combining (2) and (3), the reward change per step is bounded by:
$|\mathcal{R}(p_{i})-\mathcal{R}(p_{i-1})|\leq\lambda\varepsilon.$ (20)
This implies that the reward improvement (or degradation) at each refinement step is at most linear in the prompt change size.
Now suppose the process satisfies a minimum improvement requirement:
$\mathcal{R}(p_{i})\geq\mathcal{R}(p_{i-1})+\delta,$ (21)
for some $\delta>0$. Then combining (4) and (5), we obtain:
$\lambda\varepsilon\geq\delta\quad\Rightarrow\quad\varepsilon\geq\delta/\lambda.$ (22)
Thus, for a reward gain of at least $\delta$, the update must induce a prompt change of at least $\delta/\lambda$. Conversely, if
$\|p_{i}-p_{i-1}\|\ll\delta/\lambda,$ (23)
then the step is too small to yield meaningful improvement, and further refinement is unlikely to be effective.
This result provides a formal justification for the early stopping rule: once consecutive prompt updates fall below a semantic change threshold, reward improvement will necessarily be bounded, indicating convergence.
Tasks and Datasets
@@ Acknowledgments
General Task Evaluation.
We select six tasks from the BBH (suzgun2022challenging) and MMLU (wang2024mmlu) datasets, respectively. BBH tasks consist of six challenging reasoning tasks that assess logical inference and problem-solving skills, including boolean expressions, disambiguation QA, formal fallacies, geometric shapes, ruin names, and sports understanding. MMLU tasks cover six subject-specific tasks designed to evaluate general knowledge across diverse fields, including college biology, college medicine, electrical engineering, high school world history, human aging, and marketing.
Domain-Specific Task Evaluation.
We include three benchmarks: C-Eval (huang2024c), GSM8K (zhang2024careful), and LSAT-AR (zhong2023agieval). C-Eval is a Chinese evaluation benchmark that covers domain-specific topics such as art studies, clinical medicine, and Urban and Rural Planner. GSM8K is a widely used mathematical reasoning dataset. LSAT-AR focuses on legal reasoning, evaluating AI performance in law-related tasks.
Dataset Split.
@@
One of the key highlights of this study is that the training data consists of only a single sample—MARS utilizes just one data point from the current task for the entire optimization process. This minimal setup is enabled by the Planner agent’s strong capacity to identify the task definition and interpret the provided example, allowing it to infer the underlying task structure and semantics effectively. Since the primary function of the Planner is to generate an optimization trajectory based on the prompt-task alignment, a single representative instance is sufficient to guide downstream agents. We further analyze the impact of different sample sizes on performance in Appendix E.
@@
Experiment Settings and Baselines
We select a powerful LLM, deepseek-V2.5-1210 (guo2025deepseek), as our primary agent for the APO tasks. Not only does deepseek-V2.5-1210 exhibit strong reasoning and generation capabilities in a variety of natural language processing tasks, but it also efficiently explores multiple angles when facing complex prompt optimization requirements, making it well-suited for adapting to different tasks and datasets in the APO process. We adopt accuracy as our primary evaluation metric to comprehensively assess the performance of different methods across various task scenarios. Table 1 presents our experimental results on 12 general task datasets, illustrating the performance of APO in diverse scenarios, while Table 2 summarizes its performance on five domain-specific datasets, underscoring the model’s versatility and stability across different fields. To further validate the generality and robustness of our method, we additionally employed another high-performance LLM, GPT-4o (achiam2023gpt), for extended comparative experiments, with the corresponding findings reported in Appendix D.
We compare MARS with three categories of baselines: original prompts, CoT prompts, and some of the latest APO methods. Specifically, (1) original prompts refer to the prompts used in the datasets, where each dataset often provides some initial guidance for the tasks. (2) To build the CoT (Zero-Shot) baseline, we add the prompt Let’s think step by step at the beginning of each task; based on this, we further include a specific example to create the CoT (Few-Shot) baseline. (3) Finally, we compare MARS with some strong baseline methods from recent years, including Automatic Prompt Engineer (APE) (zhoularge), Prompt Optimization with Textual Gradients (ProTeGi) (pryzant2023automatic), Optimization by PROmpting (OPRO) (Yang0LLLZC24), and Prompt Engineer 2 (PE2) (ye2023prompt). APE and ProTeGi generate multiple prompts and perform search optimization to find the optimal prompt, while OPRO and PE2 optimize prompts by designing a sophisticated meta prompts.
Generalization Across Different Base and Target Models
@@
Base Model of GPT-4o
To verify the generality and effectiveness of the proposed method in this study, we conduct further experiments by replacing the base model with GPT-4o (achiam2023gpt).
@@
Different Initial Prompts $p_{0}$
As shown in the Table 7, MARS consistently achieves strong performance across different initial prompts $p_{0}$, demonstrating robust optimization capabilities. Although the choice of $p_{0}$ can lead to variations in absolute performance, MARS maintains a relatively stable improvement margin across tasks. This indicates that MARS effectively adapts its optimization trajectory regardless of the quality of the starting prompt, highlighting its reliability and generalization ability in diverse initialization scenarios.
@@
Model 1 2 3 Avg
$p_{0}$ Let’s think step by step. Let’s work this out step by step. Let’s proceed with our tasks one by one.
MARS 79.59 78.53 75.30 77.81
Table 7: Performance comparison of MARS on BBH tasks under different initial prompts $p_{0}$.
Sample Size Analysis
To analyze the rationality of one-shot training, we present a comparison of 0-shot, 1-shot, 3-shot, and baseline methods in Table 8.
Tasks Train B.E. D.QA $F.$ G.S. R.N. S.U.
@@
Table 8: Performance comparison of different sampling strategies on the evaluation metric. Train means the training data.
@@
Prompts for Agents
Table 9 summarizes the prompts used for all agents in this paper, each playing a crucial role in the overall optimization workflow.
The Planner first constructs a structured plan based on the task requirements, defining a trajectory of sub-goals to guide the optimization process. Beyond laying out the overall flow, the Planner provides semantic anchors that structure the interactions among downstream agents.
In the refinement phase, the Teacher generates Socratic-style questions aligned with the current sub-goal, designed to elicit reasoning rather than direct edits. The Student responds by proposing refined prompts, while the Critic evaluates the quality and pedagogical alignment of the guidance, forming an interactive loop for iterative improvement.
Finally, the Target agent validates the final prompt on downstream tasks, providing external performance feedback that closes the optimization loop. This validation ensures that the generated prompt is not only structurally coherent but also effective for the intended task.
Full-process Prompt Optimization
Figure 7 presents a comprehensive example of full-process prompt optimization, using the Geometry Shapes task from the BBH dataset. This visualization clearly illustrates the end-to-end workflow of the MARS framework, highlighting how the multi-agent system approximates a policy-guided trajectory over a discrete prompt space.
The process begins with the Planner agent, which decomposes the input task into a series of interpretable sub-goals $\{st_{1},\dots,st_{n}\}$, forming a high-level optimization trajectory. These steps serve as a form of global guidance for the subsequent reasoning path, representing the initial policy direction in the POMDP formulation.
Then, through the iterative interaction among the Teacher, Critic, and Student agents, MARS executes a sequence of transitions $(s_{i-1},a_{i},s_{i})$, where each composite action $a_{i}=(q_{i},c_{i})$ is derived from the Socratic-style question and its critique. The Student agent updates its internal latent state $s_{i}$ based on this feedback and outputs the observable prompt $o_{i}=p_{i}$. This dialogue-driven process simulates a soft pseudo-gradient trajectory in the POMDP landscape, gradually refining the prompt through interpretable and feedback-aligned steps.
MARS enables dual-level interpretability: process interpretability arises from the explicit optimization path planned by the Planner and the Socratic dialogue structure, which makes each state transition traceable and rational. Result interpretability is embodied in the final prompt, which integrates task-specific constraints—such as tolerance thresholds or validation rules—as shown in Figure 7, indicating the policy-converged output under the POMDP framework.
Universal Optimum Solution